Organic hydrogenation reactions, in which H2 and an "unsaturated" organic compound combine, are used in the food, fuel, and polymer industries. In the simplest case, ethene (C2H4) and H2 form ethane (C2H6). If 161 kJ is given off per mole of C2H4 reacting, how much heat is released when 15.9 kg of C2H6 forms?
kJ
4sf
Answer #1

C2H4 + H2 = C2H6
molar mass of C2H6 = 30.07 g/mol
mass of C2H6 = 15.9 kg = 15900 g
moles of C2H6 = mass of C2H6 / molar mass of C2H6
= 15900/30.07 = 528.8 mol
from the equation above, we know that:
moles of C2H4 = moles of C2H6 = 528.8 mol
heat produced by 528.8 mol C2H4 = 161 kJ/mol x 528.8 mol = 8.514 x 104 kJ